OpenStudy (anonymous):

A survey was undertaken at Aphane High School to determine the proportion of learners that had Facebook and Openstudy accounts. A random sample showed the following

OpenStudy (anonymous):

 If a learner had a OS account then they also had a Facebook account.  If a learner didn't have a OS account, then there was a 90% chance they had a Facebook account.  The probability that they had a OS account was 0,36

OpenStudy (anonymous):

1)Draw a Venn diagram depicting the information given above. 2)Calculate a)The probability of having a Facebook account. b) The probability of having neither a Facebook account nor a Openstudy account

OpenStudy (anonymous):

probability of having a Facebook account=0.36*1+(1-0.36)*0.9 =0.936 The probability of having neither a Facebook account nor a Openstudy account =1-0.936 =0.064

OpenStudy (anonymous):